# 11footnotetext: Equal Contribution

### v1

@@ 8 Related Works
The benefits of mismatching the training and test distribution has already been in studied in some settings. González and Abu-Mostafa (2015) demonstrate positive distribution shift in an entirely different setting, namely linear regression problems with generic mismatched training and test distributions. Unlike in our paper, they do not restrict to changing the train distribution only through data mixing, and generally only show the existence of positive distribution shift in linear rergression problems. They are able to charactarize the optimal shift explicitly only in very special cases. Canatar et al. (2021) show how to numerically optimize the training distribution in high-dimensional kernel regression problems. However, they do not characterize the positive distribution shift, but rather only show how to numerically find it for kernel regression. They do not restrict the test distribution to one coming from a data mixture. Since our focus is on mixing proportions, neither of these investigations fit our framework.
@@
There are a number of empirical papers on finding the optimal data mixture ratios. Xie et al. (2023a) and Liu et al. (2025) train a smaller proxy model to find good mixing proportions and use those for training a large model. Albalak et al. (2023) develops a multi-arm bandit algorithm for online optimization of mixing proportions. Ye et al. (2025) and Shukor et al. (2025) propose new scaling laws that depend on the mixture coefficients for determining the optimal mixing ratio for a target domain. Jiang et al. (2025)adaptively select the mixing proportions based on the scaling laws for each domain separately. The papers consider both optimizing mixing proportions so as to optimize test performance on a different target distribution, and in some papers (Xie et al., 2023a; Ye et al., 2025; Albalak et al., 2023; Jiang et al., 2025; Shukor et al., 2025) also to optimize test performance on a target which is a mixture of the training components with some fixed target proportions—as in our setup. All the papers focus on methods for empirically optimizing training proportions, rather than understanding and charactarizing the phenomena. More importnatly, in the settings considered in these papers there is significant transfer between components, which no doubts dominates the positive distribution shift—we consider a ‘pure’ setting with orthogonal tasks and no transfer to emphasize how Positive Distribution Shift can occur even in such a setting.
@@ A.1 General Power Law Tasks
For Power Law 3.1, let $f_{k}({\bm{q}})$ be approximate subpopulation error function defined as
@@
First, we show that for Power Law 3.1 and large number of samples $N$, it is sufficient to optimize over the approximate expected population loss to find $\bm{q}$}* up to error of the order $\frac{1}{N}$.
@@
For the approximate error function $f_{k}({\bm{q}})$ in Definition A.1, we have that when $Nq_{k}$ is large enough,
@@
Let $\tilde{{\bm{q}}}^{*}$ be the minimum of the approximate population error $\tilde{L}({\bm{q}})$ in Equation 5 and let $\bm{q}$$^{*}$ be the minimum of the loss in Power Law 3.1. Then if $N\geq N_{1}(p_{i},A_{i},B_{i},\alpha_{i})$
@@
Proof of Proposition A.4.
Note that by Proposition A.2, we have that for $\tilde{{\bm{q}}}^{*}$ defined in Equation 11
@@
where $C_{L}=320\min\{\frac{A_{k}}{B_{k}},1\}+2\alpha_{k}A_{k}$. Note additionally that by analogous logic from Proposition A.2 the inequality also holds the other way. By Proposition A.2, we have that
@@
Follows directly from Proposition A.4 and Proposition A.3.
@@
So it suffices to have the $o(1)$ term be smaller than $\left(\frac{p}{1-p}\right)^{\frac{1}{\alpha+1}}K^{-\frac{\alpha}{\alpha+1}}$. Note that from the proof of Proposition A.4, we can compute the $o(1)$ term. In $q_{i}^{*}$, the term was bounded by $\frac{2^{2\alpha_{1}+1}C_{L}}{C_{\lambda}N^{\frac{1}{4}}}$ where $C_{L}$ and $C_{\lambda}$ can be written explpicitly in terms of $A,B,\alpha,K,p$. In $N^{\textrm{ratio}}$ the consants are additionally mutliplied by $\frac{1}{\alpha}\left(\frac{\bigl(p^{\frac{1}{\alpha+1}}+(K-1)\bigl(\tfrac{1-p}{K-1}\bigr)^{\frac{1}{\alpha+1}}\bigr)^{\alpha+1}}{p^{1-\alpha}+(K-1)^{\alpha}(1-p)^{1-\alpha}}\right)$, so it suffices to have
@@ A.2 Memorization Tasks
Follows from Lemma A.6 and Lemma A.8. ∎
Proof Corollary 4.3.
First, note that in this case $\left(\frac{p_{K}}{p_{k}}\right)^{\frac{1}{N-1}}=\Theta((\frac{k}{K})^{\alpha/(N-1)})$. Therefore, only for $l=\Theta(K)$ do we have $f_{N}(l)=\Theta(1)$, so indeed then $K_{N}=\Theta(K)$ in this case. We directly compute that $L^{\mathrm{same}}({\bm{p}})=\Theta(N^{-1+\frac{1}{\alpha}})$. $L^{*}({\bm{p}})$ follows directly from Lemma A.8 by using $K_{N}=\Theta(K)$, and we get $L^{*}({\bm{p}})=\Theta(N^{\alpha-1})$
@@
Since $\sum_{k=1}^{K}q^{*}_{k}(N)=1$ and $q_{k}^{*}(N)=0$ for all $k>K_{N}$, by Lemma A.6, we have
@@ B.1 Proof of Main Theorem
Let $\Omega_{i}$ denote the set of ${\bm{p}}\in\Delta_{+}^{m-1}$ for which the gradient condition (14) holds for index $i\in[m]$. By Lemma B.1, the function $f_{k}$ is a rational function of $\bm{p}$}. It follows that both $\frac{\partial f_{k}({\bm{p}})}{\partial p_{i}}$ and $\sum_{k=1}^{m}p_{k}\frac{\partial f_{k}({\bm{p}})}{\partial p_{i}}$ are also rational functions of $\bm{p}$}. Therefore, $\Omega_{i}$ is the zero set of a rational function, and must be either a measure-zero subset of $\Delta_{+}^{m-1}$ or the entire domain.
Let $\Omega:=\bigcap_{i\in[m]}\Omega_{i}$ be the intersection of all $\Omega_{i}$. Then $\Omega$ is either a zero-measure subset of $\Delta_{+}^{m-1}$ or the entire domain. If $\Omega$ is a zero-measure subset, then by Lemma B.2, we are in the first case of the theorem. If $\Omega$ is the entire domain, then the gradient condition (14) holds for all $i\in[m]$, ${\bm{p}}\in\mathbb{R}^{m}_{\geq 0}\setminus\{{\bm{0}}\}$.
@@ B.2 Characterization of Conservation Conditions
Proof of Lemma 7.3.
If 7.1 holds, then for all $i,j\in[m]$ ($i\neq j$),
@@
Further, we show that if the Conservation 7.1 is satisfied, then one function $f_{i}$ determines the rest up to a constant.
@@
If both $(f_{1},\dots,f_{K},L^{\mathrm{same}})$ and $(\hat{f}_{1},\dots,\hat{f}_{K},\hat{L}^{\mathrm{same}})$ satisfy 7.1, and if $f_{i}=\hat{f}_{i}$ for some $i\in[m]$, then for all $k\neq i$, $f_{k}({\bm{p}})=\hat{f}_{k}({\bm{p}})+C_{k}$ for some constant $C_{k}$.
The above Lemma B.4 implies that for every $k$ and corresponding error function $e_{k}({\bm{n}})$, there exists at most one tuple of error functions $\{e_{j}\}_{j=1,j\neq k}^{K}$ (up to a individual constant offset for each error function $e_{j}$) that positive distribution shift does not happen for $\bm{p}$} of positive measure. This further implies the following corollary.
@@
Corollary B.5 shows that either the test mixing ratio $\bm{p}$} is on a set of measure zero on the simplex or the configuration of subpopulation error functions $e_{k}({\bm{n}})$ is on a set of measure zero. This implies that positive distribution shift exists almost always.
Proof of Lemma B.4.
Let $\Delta({\bm{p}})=L^{\mathrm{same}}({\bm{p}})-\hat{L}^{\mathrm{same}}({\bm{p}})$ be the difference between the two losses when training on the same distribution. By 7.1, we have
@@
Finally, by 7.1, we can compute for all $k\neq i$ that
@@
Proof of Corollary B.5.
This follows from Lemma B.4. Note that Lemma B.4 implies that for every $k$ and corresponding error function $e_{k}({\bm{n}})$, there exists at most one tuple of error functions $\{e_{j}\}_{j=1,j\neq k}^{K}$ (up to a individual constant offset for each error function $e_{j}$) that positive distribution shift does not happen for $\bm{p}$} of positive measure. This implies the corollary. ∎

### v2

@@ 8 Related Works
The benefits of mismatching the training and test distribution has already been in studied in some settings. gonzalez2015mismatched demonstrate positive distribution shift in an entirely different setting, namely linear regression problems with generic mismatched training and test distributions. Unlike in our paper, they do not restrict to changing the train distribution only through data mixing, and generally only show the existence of positive distribution shift in linear rergression problems. They are able to charactarize the optimal shift explicitly only in very special cases. canatar2021out show how to numerically optimize the training distribution in high-dimensional kernel regression problems. However, they do not characterize the positive distribution shift, but rather only show how to numerically find it for kernel regression. They do not restrict the test distribution to one coming from a data mixture. Since our focus is on mixing proportions, neither of these investigations fit our framework.
@@
There are a number of empirical papers on finding the optimal data mixture ratios. xie2023doremioptimizingdatamixtures and liu2025regmix train a smaller proxy model to find good mixing proportions and use those for training a large model. albalak2023efficientonlinedatamixing develops a multi-arm bandit algorithm for online optimization of mixing proportions. ye2025mixing and shukor2025scalinglawmixing propose new scaling laws that depend on the mixture coefficients for determining the optimal mixing ratio for a target domain. jiang2025adaptiveadaptively select the mixing proportions based on the scaling laws for each domain separately. The papers consider both optimizing mixing proportions so as to optimize test performance on a different target distribution, and in some papers (xie2023doremioptimizingdatamixtures; ye2025mixing; albalak2023efficientonlinedatamixing; jiang2025adaptive; shukor2025scalinglawmixing) also to optimize test performance on a target which is a mixture of the training components with some fixed target proportions—as in our setup. All the papers focus on methods for empirically optimizing training proportions, rather than understanding and charactarizing the phenomena. More importnatly, in the settings considered in these papers there is significant transfer between components, which no doubts dominates the positive distribution shift—we consider a ‘pure’ setting with orthogonal tasks and no transfer to emphasize how Positive Distribution Shift can occur even in such a setting.
@@ A.1 General Power Law Tasks
For Power Law Section 3, let $f_{k}({\bm{q}})$ be approximate subpopulation error function defined as
@@
First, we show that for Power Law Section 3 and large number of samples $N$, it is sufficient to optimize over the approximate expected population loss to find $\bm{q}$}* up to error of the order $\frac{1}{N}$.
@@
For the approximate error function $f_{k}({\bm{q}})$ in Section A.1, we have that when $Nq_{k}$ is large enough,
@@
Let $\tilde{{\bm{q}}}^{*}$ be the minimum of the approximate population error $\tilde{L}({\bm{q}})$ in Equation 5 and let $\bm{q}$$^{*}$ be the minimum of the loss in Power Law Section 3. Then if $N\geq N_{1}(p_{i},A_{i},B_{i},\alpha_{i})$
@@
Proof of Section A.1.
Note that by Section A.1, we have that for $\tilde{{\bm{q}}}^{*}$ defined in Equation 11
@@
where $C_{L}=320\min\{\frac{A_{k}}{B_{k}},1\}+2\alpha_{k}A_{k}$. Note additionally that by analogous logic from Section A.1 the inequality also holds the other way. By Section A.1, we have that
@@
Follows directly from Section A.1 and Section A.1.
@@
So it suffices to have the $o(1)$ term be smaller than $\left(\frac{p}{1-p}\right)^{\frac{1}{\alpha+1}}K^{-\frac{\alpha}{\alpha+1}}$. Note that from the proof of Section A.1, we can compute the $o(1)$ term. In $q_{i}^{*}$, the term was bounded by $\frac{2^{2\alpha_{1}+1}C_{L}}{C_{\lambda}N^{\frac{1}{4}}}$ where $C_{L}$ and $C_{\lambda}$ can be written explpicitly in terms of $A,B,\alpha,K,p$. In $N^{\textrm{ratio}}$ the consants are additionally mutliplied by $\frac{1}{\alpha}\left(\frac{\bigl(p^{\frac{1}{\alpha+1}}+(K-1)\bigl(\tfrac{1-p}{K-1}\bigr)^{\frac{1}{\alpha+1}}\bigr)^{\alpha+1}}{p^{1-\alpha}+(K-1)^{\alpha}(1-p)^{1-\alpha}}\right)$, so it suffices to have
@@ A.2 Memorization Tasks
Follows from Section A.2 and Section A.2. ∎
Proof Section 4.
First, note that in this case $\left(\frac{p_{K}}{p_{k}}\right)^{\frac{1}{N-1}}=\Theta((\frac{k}{K})^{\alpha/(N-1)})$. Therefore, only for $l=\Theta(K)$ do we have $f_{N}(l)=\Theta(1)$, so indeed then $K_{N}=\Theta(K)$ in this case. We directly compute that $L^{\mathrm{same}}({\bm{p}})=\Theta(N^{-1+\frac{1}{\alpha}})$. $L^{*}({\bm{p}})$ follows directly from Section A.2 by using $K_{N}=\Theta(K)$, and we get $L^{*}({\bm{p}})=\Theta(N^{\alpha-1})$
@@
Since $\sum_{k=1}^{K}q^{*}_{k}(N)=1$ and $q_{k}^{*}(N)=0$ for all $k>K_{N}$, by Section A.2, we have
@@ B.1 Proof of Main Theorem
Let $\Omega_{i}$ denote the set of ${\bm{p}}\in\Delta_{+}^{m-1}$ for which the gradient condition (14) holds for index $i\in[m]$. By Section B.1, the function $f_{k}$ is a rational function of $\bm{p}$}. It follows that both $\frac{\partial f_{k}({\bm{p}})}{\partial p_{i}}$ and $\sum_{k=1}^{m}p_{k}\frac{\partial f_{k}({\bm{p}})}{\partial p_{i}}$ are also rational functions of $\bm{p}$}. Therefore, $\Omega_{i}$ is the zero set of a rational function, and must be either a measure-zero subset of $\Delta_{+}^{m-1}$ or the entire domain.
Let $\Omega:=\bigcap_{i\in[m]}\Omega_{i}$ be the intersection of all $\Omega_{i}$. Then $\Omega$ is either a zero-measure subset of $\Delta_{+}^{m-1}$ or the entire domain. If $\Omega$ is a zero-measure subset, then by Section B.1, we are in the first case of the theorem. If $\Omega$ is the entire domain, then the gradient condition (14) holds for all $i\in[m]$, ${\bm{p}}\in\mathbb{R}^{m}_{\geq 0}\setminus\{{\bm{0}}\}$.
@@ B.2 Characterization of Conservation Conditions
Proof of Section 7.
If Section 7 holds, then for all $i,j\in[m]$ ($i\neq j$),
@@
Further, we show that if the Conservation Section 7 is satisfied, then one function $f_{i}$ determines the rest up to a constant.
@@
If both $(f_{1},\dots,f_{K},L^{\mathrm{same}})$ and $(\hat{f}_{1},\dots,\hat{f}_{K},\hat{L}^{\mathrm{same}})$ satisfy Section 7, and if $f_{i}=\hat{f}_{i}$ for some $i\in[m]$, then for all $k\neq i$, $f_{k}({\bm{p}})=\hat{f}_{k}({\bm{p}})+C_{k}$ for some constant $C_{k}$.
The above Section B.2 implies that for every $k$ and corresponding error function $e_{k}({\bm{n}})$, there exists at most one tuple of error functions $\{e_{j}\}_{j=1,j\neq k}^{K}$ (up to a individual constant offset for each error function $e_{j}$) that positive distribution shift does not happen for $\bm{p}$} of positive measure. This further implies the following corollary.
@@
Section B.2 shows that either the test mixing ratio $\bm{p}$} is on a set of measure zero on the simplex or the configuration of subpopulation error functions $e_{k}({\bm{n}})$ is on a set of measure zero. This implies that positive distribution shift exists almost always.
Proof of Section B.2.
Let $\Delta({\bm{p}})=L^{\mathrm{same}}({\bm{p}})-\hat{L}^{\mathrm{same}}({\bm{p}})$ be the difference between the two losses when training on the same distribution. By Section 7, we have
@@
Finally, by Section 7, we can compute for all $k\neq i$ that
@@
Proof of Section B.2.
This follows from Section B.2. Note that Section B.2 implies that for every $k$ and corresponding error function $e_{k}({\bm{n}})$, there exists at most one tuple of error functions $\{e_{j}\}_{j=1,j\neq k}^{K}$ (up to a individual constant offset for each error function $e_{j}$) that positive distribution shift does not happen for $\bm{p}$} of positive measure. This implies the corollary. ∎